OpenStudy (shelby1290):

A quadratic function is given below: f(x) = 3x^2 - x + 6 What is f(2)? a) 40 b) 28 c) 16 d) 4

OpenStudy (anonymous):

plug in x=2 wherever you see x evaluate

OpenStudy (shelby1290):

I got 16

OpenStudy (shelby1290):

Is that right?

OpenStudy (anonymous):

\(3*(2^2)-2+6\\ (3*4)-2+6\\12-2+6=10+6=16\)

OpenStudy (anonymous):

yeah

OpenStudy (shelby1290):

thanks